The weight of packets of sugar produced by a machine have a mean of 1kg and a standard deviation of 0.1kg. what is the probability that in a random sample of 50 packets the combined weight will exceed 52kg

Grade:12

1 Answers

Fanimi Olaoluwa
2 years ago
Pr( Z > 52kg)  = Pr(Z > 52 - 50(1)/50(0.1)
Pr(Z> 0.4) = 1 - Pr(Z ≤ 0)
                  = -Pr(0≤Z≤0.4)
                  = 0.5 - 0.1554
                  =0.3446